# Local Dense Logit Relations for Enhanced Knowledge Distillation

### v1

@@ 1 Introduction
Figure 1:  (a) Hinton et al. [18] introduce KD through global softmax, which calculates the probability, leading to information redundancy between classes and diminishing the logit discrimination. In classical KD, the prediction probability difference between “cat” and “dog” derived from teacher’s logit output is calculated as: $\Delta P_{KD}=|p_{1}^{t}-p_{2}^{t}|$=$|\frac{\exp({\mathbf{Z_{1}^{t}}})-\exp({\mathbf{Z_{2}^{t}}})}{\sum_{i=1}^{C}{\exp({\mathbf{Z_{i}^{t}}})}}|$. (b) In contrast, our proposed LDRLD uses category pairs and calculates probability difference between “cat” and “dog” as: $\Delta P_{LDRLD}=|p_{1}^{t}-p_{2}^{t}|$=$|\frac{\exp({\mathbf{Z_{1}^{t}}})-\exp({\mathbf{Z_{2}^{t}}})}{\sum_{i=1}^{2}{\exp({\mathbf{Z_{i}^{t}}})}}|$. It is obvious that $\Delta P_{LDRLD}>\Delta P_{KD}$, indicating that our approach enhances inter-class differences compared to KD and improves fine-grained logit discrimination.
@@ 3.2 Local Dense Relational Logit Distillation
$\mathbf{M}_{\pi(d)}=\begin{cases}-\infty&\text{ if }\pi(d)=i_{\max},\\ 1&\text{ otherwise, }\end{cases}$ (3)
@@
$\left\{\begin{array}[]{l}p_{i}^{t}=\exp\left(\mathbf{Z}_{i}^{t}/\tau\right)/S_{ij}^{t},p_{j}^{t}=\exp\left(\mathbf{Z}_{j}^{t}/\tau\right)/S_{ij}^{t},\vspace{5pt}\\ p_{i}^{s}=\exp\left(\mathbf{Z}_{i}^{s}/\tau\right)/S_{ij}^{s},p_{j}^{s}=\exp\left(\mathbf{Z}_{j}^{s}/\tau\right)/S_{ij}^{s},\\ \end{array}\right.$ (12)
@@
[…]
@@
Adaptive Decay Weight strategy. So far, we have uniformly assigned weights to all category pairs. However, this may hinder the optimization of visually complex or critical categories and lead to network performance degradation. To address this, we propose an Adaptive Decay Weight (ADW) strategy based on human perception of category similarity. The ADW integrates Inverse Rank Weighting and Exponential Rank Decay. It dynamically adjusts weights to enhance discrimination by assigning greater weights to category pairs that are more difficult for the model to distinguish (e.g., $[\mathbf{Z}_{1}^{t},\mathbf{Z}_{2}^{t}]$), and smaller weights to less related category pairs (e.g., $[\mathbf{Z}_{1}^{t},\mathbf{Z}_{4}^{t}]$). This targeted weighting sharpens the student’s focus on key categories while reducing attention to others. Details of the two key components of the ADW strategy are described below:
@@
where $Inv(x)$ denotes the inverse of $x$. $R^{\prime}$ and $R$ are the rankings of categories $i$ and $j$ in the pair $[\mathbf{Z}_{i},\mathbf{Z}_{j}]$, respectively. The term $\left|R-R^{\prime}\right|$ represents the difference in category rankings, and $\epsilon=1.50$ prevents division by zero.
@@
(2) Exponential Rank Decay. When calculating the pairwise weights according to Eq. (15), note that the pairs $[\mathbf{Z}_{1}^{t},\mathbf{Z}_{2}^{t}]$ and $[\mathbf{Z}_{12}^{t},\mathbf{Z}_{13}^{t}]$ receive the same weight. However, intuitively, distinguishing between $[\mathbf{Z}_{1}^{t},\mathbf{Z}_{2}^{t}]$ is more difficult due to highly semantic similarity, so the weight for the former should be greater than that for the latter. Based on this understanding, we propose an Exponential Ranking Decay (ERD) that dynamically adjusts weight decay according to the sum of the rankings, defined as:
@@
[…]
@@
Remaining Non-Target Knowledge. Inspired by the effectiveness of non-target knowledge in previous studies [6, 84, 76, 31], we leverage the remaining non-target classes to ensure comprehensive knowledge transfer and improve student performance. We denote the remaining logits for the teacher and student as ¯$\bar{\mathcal{H}}^{t}=[\mathbf{Z}_{d+1}^{t},\mathbf{Z}_{d+2}^{t},\ldots,\mathbf{Z}_{C}^{t}]$, and ¯$\bar{\mathcal{H}}^{s}=[\mathbf{Z}_{d+1}^{s},\mathbf{Z}_{d+2}^{s},\ldots,\mathbf{Z}_{C}^{s}]$, respectively, in Fig. 2(c). The knowledge is distilled using $\mathcal{L}_{RNTK}$, defined as follows:
@@ 4 Experiments
Distillation Teacher Publication ResNet32$\times$4 WRN40-2 ResNet50 VGG13 ResNet32$\times$4 ResNet50
@@ 1 Optimization Objective
1:$\mathcal{D}$: training dataset; $T_{\text{net}}$: pre-trained teacher network; $S_{\text{net}}$: student network with parameters $\theta$; $\eta$: learning rate
@@
7:             $\mathbf{Z}^{t}\leftarrow T_{\text{net}}(\mathbf{x}_{i})$triangleright Compute teacher’s output for sample $i$ and teacher’s parameters are fixed.
8:             $\mathbf{Z}^{s}\leftarrow S_{\text{net}}(\mathbf{x}_{i})$triangleright Compute student’s output for sample $i$.
@@ 2.1 Experimental Setups
CIFAR-100 [28] comprises 100 classes, each image with a resolution of 32$\times$32 pixels. The dataset contains 50,000 training images and 10,000 validation images.
@@
Tiny-ImageNet [30] is a streamlined version of the ImageNet-1K dataset. The dataset includes 200 classes, with images of 64$\times$64 pixels resolution. It comprises 100,000 training images and 10,000 validation images.
@@
In this work, we compare the proposed method with classical benchmarks across several datasets, including CIFAR-100, ImageNet-1K, Market-1501, Tiny-ImageNet, and COCO2017. The comparison focuses on two main categories of KD methods. We evaluate the following feature-based KD methods: FitNets [51], SP [63], CC [46], VID [1], AT [27], OFD [17], PKT [45], NST [23], RKD [44], FT [25],CRD [62], SAKD [58], ReviewKD [3], NORM [38],FCFD [37], DiffKD [22], and CAT-KD [11]. We also evaluate the following logit-based KD methods such as: KD [18], DKD [84], NKD [76], CTKD [32], LCKA [91], WTTM [87], IPWD [43], SDD [66], and LSKD [60].
@@
  • (1) For the CIFAR-100 dataset, we use a batch size of 64. The learning rate is set to 0.05 for all architectures except MobileNetV2 [52] and ShuffleNet [39, 82], for which it is set to 0.01. The hyperparameters are provided in Table 7 and Table 8, where the recursion depth is set to $d=9$ for ResNet32$\times$4 vs. ShuffleNetV1 and to $d=7$ for all other models. All models are trained with a linear warm-up for 20 epochs. See CRD [62] for detailed training details.
@@ 3.1 Hyperparameter Sensitivity
Table 12: Impact of $\alpha$ and $\beta$ on student’s performance(%) on CIFAR-100 using ResNet32$\times$4 as teacher and ShuffleNetV2 as student.
@@ 3.2 Ablation and explanation of ADW
$\mathcal{L}_{Local}$ $\mathcal{L}_{RNTK}$ ResNet32$\times$4 WRN-40-2 ResNet32$\times$4 WRN-40-2
ResNet8$\times$4 WRN-40-1 ShuffleNetV2 ShuffleNetV1
@@ 3.3 Training Efficiency
Table 16: We assess the average training time (per epoch) and test accuracy (%) using a GeForce RTX 3090 on the CIFAR-100 dataset for ResNet32$\times$4 (teacher) and ResNet8$\times$4 (student).
@@ 4.1 Differences between NCKD and RNTK
DKD decoupling results in NCKD [84], which excludes only the target class, whereas RNTK excludes the top-1 to top-$d$ most confident classes. The latter strategy avoids suppressing high-confidence categories, thereby reducing dependence on common categories, improving the ability to recognize unseen data, and promoting the activation of low-confidence classes. The experimental results demonstrate that the student with RNTK achieves the best performance when $d$=7, as shown in Table 17, outperforming NCKD and supporting this argument.
@@ 4.2 Influence of incorrect teacher predictions
Teacher ResNet110 ResNet110 ResNet32$\times$4 VGG13
Student ResNet20 ResNet32 ResNet8$\times$4 VGG8
@@ 4.4 Combination with other KD
Teacher ResNet32$\times$4 WRN-40-2 ResNet32$\times$4 VGG13
Student ResNet8$\times$4 WRN-40-1 ShuffleNetV2 VGG8
@@ 5.2 Fine-grained Task Results on the CUB200
The fine-grained task hinges on local features (e.g., feather patterns). LDRLD enhances inter-class separability by decoupling logits into combination pairs $[z_{1},z_{2}]$, amplifying subtle differences while normalizing them to reduce interference from other categories. The results in Table 22 show that LDRLD outperforms other KD and validates generalization ability (see the SDD training detail).
@@
Figure 5: Visualization of the difference in correlation matrices between student and teacher logits for different teacher-student pairs: ResNet32$\times$4 vs ResNet8$\times$4, and ResNet32$\times$4 vs ShuffleNetV2, on the CIFAR-100 dataset.

### v2

@@ 1 Introduction
Figure 1:  (a) Hinton et al. [18] introduce KD through global softmax, which calculates the probability, leading to information redundancy between classes and diminishing the logit discrimination. In classical KD, the prediction probability difference between “cat” and “dog” derived from teacher’s logit output is calculated as: $\Delta P_{KD}=|p_{1}^{t}-p_{2}^{t}|$$|\frac{\exp({\mathbf{Z_{1}^{t}}})-\exp({\mathbf{Z_{2}^{t}}})}{\sum_{i=1}^{C}{\exp({\mathbf{Z_{i}^{t}}})}}|$. (b) In contrast, our proposed LDRLD uses category pairs and calculates probability difference between “cat” and “dog” as: $\Delta P_{LDRLD}=|p_{1}^{t}-p_{2}^{t}|$$|\frac{\exp({\mathbf{Z_{1}^{t}}})-\exp({\mathbf{Z_{2}^{t}}})}{\sum_{i=1}^{2}{\exp({\mathbf{Z_{i}^{t}}})}}|$. It is obvious that $\Delta P_{LDRLD}>\Delta P_{KD}$, indicating that our approach enhances inter-class differences compared to KD and improves fine-grained logit discrimination.
@@ 3.2 Local Dense Relational Logit Distillation
[…]
@@
[…]
@@
$\begin{aligned} \mathcal{L}&=\sum_{i=1}^{d-1}\mathcal{KL}\left(\mathfrak{C}_{i}^{t},\mathfrak{C}_{i}^{s}\right)\\ &=\sum_{i=1}^{d-1}\sum_{j=i+1}^{d}\mathcal{KL}\left(\sigma(\phi(\mathbf{Z}_{i}^{t},\mathbf{Z}_{j}^{t}),\tau),\sigma(\phi(\mathbf{Z}_{i}^{s},\mathbf{Z}_{j}^{s}),\tau)\right)\\ &=\sum_{i=1}^{d-1}\sum_{j=i+1}^{d}\left[p_{i}^{t}\log(p_{i}^{t}/p_{i}^{s})+p_{j}^{t}\log(p_{j}^{t}/p_{j}^{s})\right].\end{aligned}$
@@
Adaptive Decay Weight strategy. So far, we have uniformly assigned weights to all category pairs. However, this may hinder the optimization of visually complex or critical categories and lead to network performance degradation. To address this, we propose an Adaptive Decay Weight (ADW) strategy based on human perception of category similarity. The ADW integrates Inverse Rank Weighting and Exponential Rank Decay. It dynamically adjusts weights to enhance discrimination by assigning greater weights to category pairs that are more difficult for the model to distinguish (e.g., $\mathbf{Z}_{1}^{t},\mathbf{Z}_{2}^{t}]$), and smaller weights to less related category pairs (e.g., $\mathbf{Z}_{1}^{t},\mathbf{Z}_{4}^{t}]$). This targeted weighting sharpens the student’s focus on key categories while reducing attention to others. Details of the two key components of the ADW strategy are described below:
@@
where $Inv(x)$ denotes the inverse of $x$. $R^{\prime}$ and $R$ are the rankings of categories $i$ and $j$ in the pair $\mathbf{Z}_{i},\mathbf{Z}_{j}]$, respectively. The term $\left|R-R^{\prime}\right|$ represents the difference in category rankings, and $\epsilon=1.50$ prevents division by zero.
@@
(2) Exponential Rank Decay. When calculating the pairwise weights according to Eq. (15), note that the pairs $\mathbf{Z}_{1}^{t},\mathbf{Z}_{2}^{t}]$ and $\mathbf{Z}_{12}^{t},\mathbf{Z}_{13}^{t}]$ receive the same weight. However, intuitively, distinguishing between $\mathbf{Z}_{1}^{t},\mathbf{Z}_{2}^{t}]$ is more difficult due to highly semantic similarity, so the weight for the former should be greater than that for the latter. Based on this understanding, we propose an Exponential Ranking Decay (ERD) that dynamically adjusts weight decay according to the sum of the rankings, defined as:
@@
$\begin{aligned} \mathcal{L}^{w}=\sum_{i=1}^{d-1}{\sum_{j=i+1}^{d}{\Omega_{ADW}}}\left(i,j\right)\left[p_{i}^{t}\log(p_{i}^{t}/p_{i}^{s})+p_{j}^{t}\log(p_{j}^{t}/p_{j}^{s})\right].\end{aligned}$
@@
Remaining Non-Target Knowledge. Inspired by the effectiveness of non-target knowledge in previous studies [6, 84, 76, 31], we leverage the remaining non-target classes to ensure comprehensive knowledge transfer and improve student performance. We denote the remaining logits for the teacher and student as $\bar{\mathcal{H}}^{t}=[\mathbf{Z}_{d+1}^{t},\mathbf{Z}_{d+2}^{t},\ldots,\mathbf{Z}_{C}^{t}]$, and $\bar{\mathcal{H}}^{s}=[\mathbf{Z}_{d+1}^{s},\mathbf{Z}_{d+2}^{s},\ldots,\mathbf{Z}_{C}^{s}]$, respectively, in Fig. 2(c). The knowledge is distilled using $\mathcal{L}_{RNTK}$, defined as follows:
@@ 4 Experiments
[…]
@@ 1 Optimization Objective
[…]
@@
7:     $\mathbf{Z}^{t}\leftarrow T_{\text{net}}(\mathbf{x}_{i})$ $\triangleright$ Compute teacher’s output for sample $i$ and teacher’s parameters are fixed.
8:     $\mathbf{Z}^{s}\leftarrow S_{\text{net}}(\mathbf{x}_{i})$ $\triangleright$ Compute student’s output for sample ii.
@@ 2.1 Experimental Setups
CIFAR-100 [28] comprises 100 classes, each image with a resolution of $\times32$ pixels. The dataset contains 50,000 training images and 10,000 validation images.
@@
Tiny-ImageNet [30] is a streamlined version of the ImageNet-1K dataset. The dataset includes 200 classes, with images of $\times64$ pixels resolution. It comprises 100,000 training images and 10,000 validation images.
@@
In this work, we compare the proposed method with classical benchmarks across several datasets, including CIFAR-100, ImageNet-1K, Market-1501, Tiny-ImageNet, and COCO2017. The comparison focuses on two main categories of KD methods. We evaluate the following feature-based KD methods: FitNets [51], SP [63], $C$ [46], VID [1], AT [27], OFD [17], PKT [45], NST [23], RKD [44], FT [25],CRD [62], SAKD [58], ReviewKD [3], NORM [38],FCFD [37], DiffKD [22], and CAT-KD [11]. We also evaluate the following logit-based KD methods such as: KD [18], DKD [84], NKD [76], CTKD [32], LCKA [91], WTTM [87], IPWD [43], SDD [66], and LSKD [60].
@@
[…]
@@ 3.1 Hyperparameter Sensitivity
[…]
@@ 3.2 Ablation and explanation of ADW
[…]
@@ 3.3 Training Efficiency
[…]
@@ 4.1 Differences between NCKD and RNTK
DKD decoupling results in NCKD [84], which excludes only the target class, whereas RNTK excludes the top-1 to top-$d$ most confident classes. The latter strategy avoids suppressing high-confidence categories, thereby reducing dependence on common categories, improving the ability to recognize unseen data, and promoting the activation of low-confidence classes. The experimental results demonstrate that the student with RNTK achieves the best performance when dd=7, as shown in Table 17, outperforming NCKD and supporting this argument.
@@ 4.2 Influence of incorrect teacher predictions
[…]
@@ 4.4 Combination with other KD
[…]
@@ 5.2 Fine-grained Task Results on the CUB200
The fine-grained task hinges on local features (e.g., feather patterns). LDRLD enhances inter-class separability by decoupling logits into combination pairs $z_{1},z_{2}]$, amplifying subtle differences while normalizing them to reduce interference from other categories. The results in Table 22 show that LDRLD outperforms other KD and validates generalization ability (see the SDD training detail).
@@
[…]